OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. x2 - 16 = 0 and x2 = -2x + 24 Part 2: Describe what the solution(s) represent to the graph of each. Part 3: How are the graphs alike? How are they different?

OpenStudy (amorfide):

x²-16 should be easily spotted as the difference of two squares so square root 16 and x so you get x+4 then you have one other set of brackets to give -16 and 0 x's difference of two squares it the same numbers different signs so for example 

OpenStudy (anonymous):

Yes thank you, I actually already had gotten the first though. The one that I was having trouble with was finding x2 = -2x + 24

OpenStudy (amorfide):

x²=-2x+24 add 2x to both sides and subtract 24 from both sides to make this equal to 0 x²+2x-24=0 now we shall factorise this multiply -24 by coefficient of x² 1x-24=-24 we need two numbers to add up to coefficient of x which is 2 and multiply to give -24 +6 and -4 work for this solution now replace 2x with +6x and -4x to get x²+6x-4x-24 take out the factor in each x(x+6)-4(x+6) (x-4)(x+6)

OpenStudy (amorfide):

Part 2: Describe what the solution(s) represent to the graph of each. (x+4)(x-4) this graph means that the graph crosses the x axes at two points x+4=0 x=-4 this cuts at -4 on the x axes x-4=0 x=4 this cuts at 4 on the x axes so it cuts x axes at 4 and -4 (x-4)(x+6) this graph cuts at x-4=0 x=4 x+6=0 x=-6 this cuts at 4 and -6 on the x axes

OpenStudy (anonymous):

Thank you so much that helped a lot!

OpenStudy (amorfide):

Part 3 they are both positive x squared so they have the same shape (positive parabola i think it is called) they both cut at 4 on the x axes they are different because the second equation cuts at -6 rather than -4